OpenStudy (anonymous):

Use the substitution method to solve the linear system given. Express the answer as an ordered pair of the form (x,y) y = x + 2 3x + 2y = 9

OpenStudy (anonymous):

Substitute y = x + 2 into 3x + 2y = 9 Making the equation 3x + 2(x+2) = 9 Slowly Expanding the equation 3x + 2x + 4 = 9 Rearranging and balancing out the equation 3x + 2x = 9 - 4 Adding like terms 5x = 5 Dividing both sides by 5 to find x: x = 1 Substitute x = 1 into first said equation (y = x + 2) y = 1 + 2 y = 3 Therefore, (x, y) is (1, 3).